OpenStudy (anonymous):

Lisa wants to buy a phone cards for $5. Company X charges $0.37 for the first minute and $0.08 for each additional minute. Company Y charges $0.52 for the first minute and $0.14 for each additional minute. If Lisa makes just one call, which phone company gives her more minutes for $5? How many more minutes? A. Company Y; about 29 minutes B. Company X; about 29 minutes C. Company X; about 26 minutes D. Company Y; about 26 minutes

OpenStudy (anonymous):

Company X: 5 = 0.37 + 0.08x Company Y: 5 = 0.52 + 0.14x

OpenStudy (anonymous):

Solving for x you find: Company X gives about 58 minutes Company Y gives about 32 minutes. Company X gives more minutes by about 26

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

No problem.